# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 11.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.183183183The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.184184184Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 11.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.185185185The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates186186186At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@
Table 22 and Figure 20 report cumulative regret at four checkpoints. The results confirm the theoretical progression from Table 21: $\varepsilon$-greedy grows linearly, UCB1 and Thompson Sampling grow as $\sqrt{T}$, and UCB-PI-tuned achieves the lowest regret at every checkpoint past T=10, [^0] T=10{,}000. The untuned UCB-PI variant overexplores, scaling as $\sqrt{T}$ rather than the theoretical $\log T$, illustrating that WARP-based elimination alone is insufficient without variance-calibrated exploration bonuses.
Table 22: Cumulative regret for six pricing algorithms on the Misra et al. (2019) demand environment ($K=100$, S=1, [^0] S=1{,}000, 10 seeds).
Level Algorithm T=10T{=}10K T=50T{=}50K T=100T{=}100K T=200T{=}200K Rate
0 $\varepsilon$-greedy 160 624 1,158 2,263 $\Theta(T)$
1 LTE (5%) 1,005 1,165 1,363 1,771 $O(T^{2/3})$
2 UCB1 719 2,609 4,263 6,734 $O(\sqrt{KT})$
3 Thompson 290 628 844 1,136 $O(\sqrt{KT})$
4 UCB-PI 789 2,266 3,244 4,503 $O(\log T)$
5 UCB-PI-tuned 338 541 640 780 $O(\log T)$
@@
Figure 20: Cumulative regret on log-log axes. Dashed lines show $\Theta(T)$, $O(\sqrt{T})$, and $O(\log T)$ reference rates. Shaded regions are $\pm 2$ standard errors across 10 seeds.
12 Reinforcement Learning from Human Feedback

### 17 Mar 2026 v3

@@ 10.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.190190190The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.191191191Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 10.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.192192192The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates193193193At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@
Figure 19 reports cumulative regret at four checkpoints. The results confirm the theoretical progression from Table 19: $\varepsilon$-greedy grows linearly, UCB1 and Thompson Sampling grow as $\sqrt{T}$, and UCB-PI-tuned achieves the lowest regret at every checkpoint past T=10, [^0] T=10{,}000. The untuned UCB-PI variant overexplores, scaling as $\sqrt{T}$ rather than the theoretical $\log T$, illustrating that WARP-based elimination alone is insufficient without variance-calibrated exploration bonuses.
@@
Figure 19: Cumulative regret on log-log axes ($K=100$, S=1, [^0] S=1{,}000, 10 seeds). Each algorithm’s legend entry includes its theoretical regret rate. Dashed lines show $\Theta(T)$, $O(\sqrt{T})$, and $O(\log T)$ reference rates. Shaded regions are $\pm 2$ standard errors.
11 Offline Reinforcement Learning and Human Feedback
In the preceding chapters, the agent interacts with its environment while learning: it tries a price, observes a purchase, and updates its belief. Online learning is natural in digital markets where experimentation is cheap and feedback is instant. In many economically important settings, however, experimentation is impossible or prohibitively costly. A hospital cannot randomly assign treatments to learn optimal dosing. A central bank cannot experiment with interest rate schedules to discover optimal monetary policy. A firm inheriting a decade of transaction logs wants to improve its pricing rule without conducting new experiments during the transition. In each case, the agent has access to a fixed dataset of past decisions and outcomes, collected under some historical policy, and must learn the best possible new policy from this data alone.
This is the problem of offline reinforcement learning, also called batch reinforcement learning.194194194The term “batch RL” was standard in the earlier literature (Ernst et al., 2005; Lange et al., 2012). “Offline RL” became dominant after Levine et al. (2020), who distinguished it from off-policy RL (which still collects new data, just under a different policy than the target). I use “offline RL” throughout. The agent never queries the environment. All learning happens from a fixed dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})\}_{i=1}^{n}$ collected by some behavioral policy $\pi_{b}$. The goal is to find a policy $\hat{\pi}$ whose value $V^{\hat{\pi}}$ is as close to the optimal $V^{*}$ as possible.
Online RL algorithms (Q-learning, SARSA, policy gradient methods from Section 4) can in principle be applied to offline data by treating the dataset as a replay buffer. In practice, this fails catastrophically. The reason is distributional shift: the learned policy $\hat{\pi}$ inevitably queries state-action pairs that the behavioral policy $\pi_{b}$ never visited, and the Q-function at these unseen pairs is pure extrapolation. Because the Bellman backup propagates these extrapolation errors through the max operator, errors compound geometrically across the planning horizon, producing arbitrarily poor policies.195195195This failure mode was first demonstrated empirically by Fujimoto et al. (2019), who showed that standard off-policy algorithms (DDPG, SAC) trained purely from a static dataset performed worse than the behavioral policy itself, even when that behavioral policy was a partially-trained, mediocre agent. The gap widened with dataset size, the opposite of what one expects from more data.
11.1 The Pessimism Principle
The overestimation failure has a clean theoretical characterization. Consider tabular Q-learning with a fixed dataset. At any state-action pair $(s,a)$ not in the dataset, the empirical Bellman backup is undefined, but the max in $\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime})$ may still select it if the randomly-initialized Q-value happens to be high. With function approximation, the problem is subtler but identical in spirit: the function approximator can assign high values to out-of-distribution inputs without any corrective signal.
The solution, established simultaneously by several groups, is the pessimism principle: construct a lower confidence bound on the Q-function and optimize against it. Formally, given a dataset $\mathcal{D}$ and a confidence parameter $\delta$, construct a penalty function $\Gamma(s,a)$ that is large where data coverage is poor, and define the pessimistic Q-function
$\tilde{Q}(s,a)=\hat{Q}(s,a)-\Gamma(s,a)$ (94)
where $\hat{Q}$ is the standard empirical Bellman solution. The policy $\hat{\pi}(s)=\arg\max_{a}\tilde{Q}(s,a)$ selects actions that are both high-value and well-supported by data.
Definition D1 (Pessimistic Value Iteration, PEVI (Jin et al., 2021)).
Given dataset $\mathcal{D}$, penalty function $\Gamma_{h}(s,a)$ for each stage $h$, and horizon $H$, PEVI computes
$\hat{Q}_{h}(s,a)$ $=r_{h}(s,a)+\hat{P}_{h}\hat{V}_{h+1}(s,a)-\Gamma_{h}(s,a)$ (95)
$\hat{V}_{h}(s)$ $=\max\{\max_{a}\hat{Q}_{h}(s,a),\,0\}$ (96)
$\hat{\pi}_{h}(s)$ $=\arg\max_{a}\hat{Q}_{h}(s,a)$ (97)
where $\hat{P}_{h}$ is the empirical transition operator estimated from $\mathcal{D}$.
Jin et al. (2021) show that with $\Gamma_{h}(s,a)=c\cdot\sqrt{H^{3}/N_{h}(s,a)}$, where $N_{h}(s,a)$ counts visits to $(s,a)$ at stage $h$ and $c$ is an absolute constant, PEVI achieves
$V^{*}-V^{\hat{\pi}}\leq\tilde{O}\left(\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}}\left[\sqrt{\frac{1}{N_{h}(s_{h},a_{h})}}\right]\right)$ (98)
with high probability. The bound depends on the data coverage at states and actions visited by the optimal policy $\pi^{*}$, not the full state-action space. This is the key advantage of pessimism over uniform coverage requirements.
11.1.1 Concentrability and Coverage
The bound in (98) is instance-dependent, scaling with how well $\pi_{b}$ covers $\pi^{*}$. The classical way to formalize this is through concentrability coefficients (Munos and Szepesvári, 2008).
Definition D2 (Single-policy concentrability).
The single-policy concentrability coefficient of $\pi^{*}$ with respect to $\pi_{b}$ is
$C^{*}=\max_{h\in[H]}\left\|\frac{d_{h}^{\pi^{*}}}{d_{h}^{\pi_{b}}}\right\|_{\infty}$ (99)
where $d_{h}^{\pi}(s,a)$ is the state-action occupancy measure of $\pi$ at stage $h$.
When $C^{*}$ is finite, the optimal policy visits only states and actions that $\pi_{b}$ also visits with non-negligible probability, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (98) remain controlled. Rashidinejad et al. (2021) prove that single-policy concentrability is both necessary and sufficient for offline learning: if $C^{*}<\infty$, then $O(|\mathcal{S}|H^{2}C^{*}/\epsilon^{2})$ samples suffice for an $\epsilon$-optimal policy; if $C^{*}=\infty$, no algorithm can guarantee suboptimality better than the behavioral policy.
11.1.2 Impossibility Results
Zanette (2021) establish fundamental limits on offline RL by showing that it can be exponentially harder than online RL. Specifically, there exist MDPs with $S$ states and horizon $H$ where online RL finds an $\epsilon$-optimal policy in $\text{poly}(S,H,1/\epsilon)$ episodes, but any offline algorithm requires $\Omega(2^{H})$ samples unless the dataset covers all reachable states. The construction uses a binary tree MDP where the optimal path visits a unique leaf, and any dataset that misses this leaf provides no information about the optimal action at the root.
The practical implication is that offline RL is not a universal replacement for online experimentation. When the behavioral policy is far from optimal, especially in long-horizon problems, offline methods provably cannot recover the optimal policy without exponentially large datasets. Pessimistic algorithms are the best one can do, but they are still fundamentally constrained by what the data contains.
11.2 Algorithms
I present four practical algorithms that instantiate different approaches to the distributional shift problem. All four can be understood as modifications of standard Q-learning (Section 4) that prevent the agent from overvaluing actions outside the data support.
11.2.1 Fitted Q-Iteration
Fitted Q-Iteration (FQI, Ernst et al., 2005) is the simplest offline RL algorithm, predating the modern pessimism framework. FQI applies the standard Bellman backup iteratively using supervised regression on the fixed dataset, as described in Section 4. It does not include any explicit pessimism mechanism. When state-action coverage is poor, the max in the Bellman backup selects the highest Q-value among all actions at $s^{\prime}$, including actions never observed in the data. If the function approximator generalizes poorly at these unseen actions, targets become noisy and biased upward, causing the overestimation cascade. FQI works well when coverage is good but degrades as coverage gaps grow.196196196Munos and Szepesvári (2008) prove finite-time error bounds for FQI under approximate Bellman completeness and all-policy concentrability. Both assumptions are strong, and violation of either leads to divergence in practice.
11.2.2 Conservative Q-Learning
Conservative Q-Learning (CQL, Kumar et al., 2020) adds an explicit penalty that pushes down Q-values at actions not well-represented in the data. The key idea is to add a regularizer to the Bellman error objective that minimizes Q-values under a broad distribution over actions while maximizing Q-values at the actions actually taken in the dataset.
Definition D3 (Conservative Q-Learning).
CQL modifies the standard Bellman error objective by adding a conservative regularizer. At each iteration, CQL solves
$\hat{Q}_{k+1}=\arg\min_{Q}\;\alpha\left(\mathbb{E}_{s\sim\mathcal{D}}\left[\log\sum_{a}\exp Q(s,a)\right]-\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]\right)+\frac{1}{2}\mathbb{E}_{\mathcal{D}}\left[(Q(s,a)-\hat{\mathcal{B}}^{\pi_{k}}\hat{Q}_{k}(s,a))^{2}\right]$ (100)
where $\alpha>0$ is a hyperparameter controlling the degree of conservatism, and $\hat{\mathcal{B}}^{\pi_{k}}$ is the empirical Bellman operator.
The first term in the regularizer, $\log\sum_{a}\exp Q(s,a)$, is a soft maximum over all actions, pushing down Q-values uniformly. The second term, $\mathbb{E}_{\mathcal{D}}[Q(s,a)]$, pulls Q-values back up at the data actions. The net effect is a penalty on Q-values for actions that appear infrequently relative to their softmax contribution. Kumar et al. (2020) prove that the resulting Q-function is a pointwise lower bound on the true Q-function (Theorem 3.2), making it a concrete instantiation of the pessimism principle from (94) with an implicit, data-adaptive penalty $\Gamma$.
11.2.3 Implicit Q-Learning
Implicit Q-Learning (IQL, Kostrikov et al., 2022) avoids querying Q-values at unseen actions entirely. Instead of computing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ in the Bellman backup (which requires evaluating $Q$ at potentially out-of-distribution actions), IQL learns a separate value function $V(s)$ that approximates the in-sample maximum through expectile regression.
Definition D4 (Implicit Q-Learning).
IQL maintains three functions: $Q_{\theta}(s,a)$, $V_{\psi}(s)$, and a policy $\pi_{\phi}(a|s)$. The value function is trained via expectile regression
$L_{V}(\psi)=\mathbb{E}_{(s,a)\sim\mathcal{D}}\left[L_{2}^{\tau}(Q_{\bar{\theta}}(s,a)-V_{\psi}(s))\right]$ (101)
where $L_{2}^{\tau}(u)=|\tau-\mathbbm{1}\{u<0\}|\cdot u^{2}$ is the asymmetric squared loss with expectile parameter $\tau\in(0.5,1)$, and $Q_{\bar{\theta}}$ uses a target network. The Q-function is trained with $V_{\psi}$ as the continuation value
$L_{Q}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(r+\gamma V_{\psi}(s^{\prime})-Q_{\theta}(s,a))^{2}\right]$ (102)
The expectile parameter $\tau$ controls the degree of optimism within the data support. As $\tau\to 1$, the expectile converges to the in-sample maximum; as $\tau\to 0.5$, it converges to the in-sample mean. Setting $\tau=0.7$ balances exploiting the best observed actions against the noise in finite samples. The critical property is that the Q-function is never evaluated at actions outside the dataset: the $\max$ operation is implicit in the expectile regression of $V$.
11.2.4 Batch-Constrained Q-Learning
Batch-Constrained Q-Learning (BCQ, Fujimoto et al., 2019) takes a different approach: rather than modifying the Q-function objective, it restricts the policy to actions similar to those in the dataset. BCQ first learns a generative model $G_{\omega}(s)$ of the behavioral policy, then constrains the policy to only select actions with high likelihood under $G_{\omega}$.
Definition D5 (Batch-Constrained Q-Learning).
BCQ learns a behavioral model $G_{\omega}(a|s)$ from the dataset via maximum likelihood, and constrains action selection
$\hat{\pi}(s)=\arg\max_{a:G_{\omega}(a|s)\geq\tau\cdot\max_{a^{\prime}}G_{\omega}(a^{\prime}|s)}Q_{\theta}(s,a)$ (103)
where $\tau\in(0,1]$ is a threshold parameter. The Q-function is trained with standard Bellman backups, but the max in the target computation is also constrained to the feasible action set.
BCQ’s constraint is hard rather than soft: actions with behavioral probability below $\tau$ times the most likely action are simply excluded from consideration. This prevents the Q-function from ever being queried at truly out-of-distribution actions, addressing the distributional shift problem at the policy level rather than the value function level. The tradeoff is that BCQ’s performance is bounded by the quality of actions in the dataset. If the behavioral policy never takes the optimal action at some state, BCQ cannot discover it regardless of sample size.
11.3 Simulation: Offline RL for Dynamic Pricing
The simulation study evaluates the four offline RL algorithms on a perishable inventory pricing problem with demand regime switching. A retailer with $I_{\max}=30$ units of perishable inventory must set prices over $H=20$ periods. The state $(i,d,t)$ consists of current inventory $i\in\{0,\ldots,30\}$, demand regime $d\in\{1,2,3,4\}$, and time remaining $t\in\{1,\ldots,20\}$. The action is a price $p\in\{1,\ldots,10\}$. Demand follows $Q\sim\text{Poisson}(\lambda_{0}[d]\cdot e^{-0.15p})$ with base rates $\lambda_{0}=(1.5,3.0,5.0,8.0)$, and the reward is $r=p\cdot\min(Q,i)$. Demand regimes follow a 4-state Markov chain with diagonal persistence 0.6. Unsold inventory at the terminal period incurs a spoilage cost of \$2.00 per unit, making clearance pricing valuable near the deadline.197197197The spoilage penalty creates distributional shift. The optimal policy adapts prices to inventory level and time remaining, using lower prices near the deadline when inventory is high. The behavioral policy ignores these state variables and prices at the maximum, so the Q-function at state-adapted pricing actions is extrapolation from sparse data. The \$2.00 penalty is a deliberate design choice: under harsher penalties (e.g., \$10 per unit), all methods collapse to 48–53% of optimal regardless of algorithmic sophistication, confirming that no offline correction can overcome severe distributional shift when the penalty regime amplifies consequences of the behavioral policy’s suboptimality. The behavioral policy represents a conservative pricing team that always sets the maximum price ($p=10$) regardless of demand regime, inventory, or time remaining, with probability 0.85 and randomizes uniformly over all prices with probability 0.15.198198198With 500 episodes (10,000 transitions) on a state-action space of 24,800 pairs, the 85% concentration at price 10 ensures that the behavioral state-action occupancy diverges significantly from the optimal policy’s occupancy, while the 15% uniform component provides sparse off-policy coverage. All episodes start at full inventory ($i=30$). All offline methods train on 500 episodes and are evaluated over 1,000 episodes against the DP optimal policy computed by backward induction.199199199FQI uses the standard Bellman backup with $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$, deliberately without a target network, to isolate the overestimation cascade as a pedagogical baseline. Adding target networks to FQI mitigates but does not eliminate extrapolation error. CQL and IQL include target networks following their original implementations (Kumar et al., 2020; Kostrikov et al., 2022); for CQL, target networks proved essential, as the conservative penalty amplifies bootstrap instability without them. Results are averaged over 20 independent seeds.
Table 20: Policy value for each offline RL method, expressed as mean return and percentage of the DP optimal. Standard errors computed over 20 seeds.
Method Mean Return % of Optimal
DP Oracle $192.41\pm 0.33$ $100.0\%$
BC $169.27\pm 0.60$ $88.0\%$
FQI $156.18\pm 1.68$ $81.2\%$
CQL $176.73\pm 1.13$ $91.9\%$
IQL $176.98\pm 0.56$ $92.0\%$
BCQ $169.27\pm 0.60$ $88.0\%$
FQI achieves 81.2% of the DP optimal, substantially below the behavioral cloning baseline of 88.0% (Table 20). Without any mechanism to control extrapolation error, the $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ operator in FQI’s Bellman backup selects overestimated Q-values at out-of-distribution actions, and these overestimates compound across 200 iterations of fitted Q-iteration. CQL and IQL both exceed the behavioral baseline, achieving 91.9% and 92.0% respectively.200200200CQL uses $\alpha=0.1$, the result of a search over $\alpha\in\{5.0,2.0,0.5,0.1\}$. Larger values push Q-values down too aggressively, collapsing the learned policy to the behavioral action at most states; the right $\alpha$ is problem-specific and can vary by orders of magnitude. CQL’s conservative penalty suppresses Q-values at actions not well-represented in the data while preserving relative ordering among data-supported actions, allowing the policy to discover lower prices that clear inventory near the deadline. IQL achieves the same improvement through a different mechanism: by replacing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ with the expectile-regressed value function $V(s^{\prime})$ as the Bellman continuation, IQL avoids querying Q at unseen actions during training while still extracting a policy that improves on the behavioral at states where the 15% noise component revealed better pricing actions. BCQ matches the behavioral at 88.0%; its action constraint restricts the policy to prices near 10, preventing both overestimation and improvement.201201201When the behavioral policy concentrates 85% probability at a single action, BCQ’s threshold constraint permits only that action at most states, effectively reducing BCQ to behavioral cloning regardless of the learned Q-values.
These results validate several theoretical predictions from the preceding sections. FQI’s degradation below the behavioral baseline (81.2% versus 88.0%) demonstrates the overestimation cascade described in Section 11.2: without pessimism, the $\max_{a^{\prime}}$ operator selects overestimated Q-values at out-of-distribution actions, and 200 iterations of bootstrapping compound these errors geometrically (Fujimoto et al., 2019). CQL and IQL both exceeding BC confirms the pessimism principle (Section 11.1): CQL’s conservative penalty produces a pointwise lower bound on the true Q-function (Definition D3, Theorem 3.2 of Kumar et al. 2020), while IQL’s expectile regression (Definition D4) avoids querying Q at unseen actions entirely (Kostrikov et al., 2022). BCQ matching BC exactly illustrates the action-constraint tradeoff formalized in Definition D5: performance is bounded by the quality of actions in the data, and when the behavioral policy concentrates on a single action, no amount of Q-learning can overcome the constraint.
Figure 20: Policy value (as % of DP optimal) versus behavioral policy randomness $\epsilon_{b}$ for four offline RL methods and the behavioral cloning baseline (BC). Higher $\epsilon_{b}$ increases data coverage. FQI peaks at moderate coverage ($\epsilon_{b}=0.3$) and collapses at both extremes. BCQ degrades at high $\epsilon_{b}$ as its behavioral constraint becomes vacuous.
Figure 20 varies the behavioral noise parameter $\epsilon_{b}\in\{0.05,0.3,0.9\}$. CQL and IQL remain above the BC baseline across all coverage levels, confirming that both pessimism mechanisms provide robustness to the data distribution. FQI exhibits non-monotone behavior: it peaks at moderate coverage ($\epsilon_{b}=0.3$) where partial exploration controls the overestimation cascade, but collapses at both extremes.202202202At high $\epsilon_{b}$, the near-uniform behavioral policy provides Q-function targets across all actions, giving the unconstrained $\max_{a^{\prime}}$ operator more opportunities to select overestimated values rather than fewer. BCQ collapses at $\epsilon_{b}=0.9$ because a nearly uniform behavioral policy renders the action constraint vacuous, reducing BCQ to unconstrained FQI. These coverage patterns connect directly to the concentrability framework (Definition D2): as $\epsilon_{b}$ decreases, the concentrability coefficient $C^{*}$ grows because the optimal policy’s state-action occupancy diverges from the behavioral policy’s, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (98) become large at states the optimal policy visits. CQL and IQL remain robust because their conservative adjustments scale with data sparsity, while FQI lacks this adaptive correction.
11.4 From Offline RL to Human Feedback
The offline RL algorithms above assume access to a scalar reward signal in the dataset. When rewards are observed, the problem reduces to learning a good policy from fixed data. In many domains, however, the reward itself is unknown and must be learned from human judgments. Reinforcement Learning from Human Feedback (RLHF) combines offline preference data with the policy optimization tools of offline RL: the agent never interacts with the environment during training, and all learning proceeds from a static dataset of human comparisons. The methods below extend the offline RL framework from learning policies given rewards to learning rewards given preferences, and then optimizing policies against those learned rewards.